To know the resistance G of a galvanometer by half deflection method, a battery of emf VE and resistance R is used to deflect the galvanometer by angle θ. If a shunt of resistance S is needed to get half deflection then G,R and S are related by the equation :
(1)2S(R+G)=RG
(2)S(R+G)=RG
(3)2S=G
(4)2G=S
1
Expert's answer
2019-11-20T09:42:19-0500

The current without the shunt:


"I_1=\frac{V_E}{R+G},"

The current with the shunt (it is parallel to the galvanometer):


"I_2=\frac{V_E}{R+\frac{GS}{G+S}},"

and this current is twice as small as the current in the first case:


"I_2=I_1\/2."

Solve this by equating the currents:


"S(R+G)+RG"